OpenStudy (lisa123):

2 4 ln(x) − ln(x + 6) − ln(x − 6)

OpenStudy (solomonzelman):

Use the following rules. \(\large\color{#000000 }{ \displaystyle c\ln(a)=\ln(a^c) }\) \(\large\color{#000000 }{ \displaystyle \ln(a\cdot b)=\ln(a)+\ln(b) }\) \(\large\color{#000000 }{ \displaystyle \ln(a\div b)=\ln(a)-\ln(b) }\) (I`m assuming you need to write everything as a single logarithm)